Find the determinant of the following matrices and indicate if any of them are singular. Show your work or use excel formulas (NOT the mdeterm function!)

a)

1/3 2/3
2 4

b)

4 -3
6 -2

c)

8 -1 3
4 2 -4
-3 1 3
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

a)

Det=4*(1/3)-2*(2/3)=0

It is a singular matrix

b)

Det=4*(-2)-(-3)*(6)=10

It is not a singular matrix

c)

Det=8*(2*3+4)+1*(4*3-4*3)+3*(4*1-(-3)*2)=110

It is not a singular matrix

Kindly revert for any queries

Thanks.